Which one of the following charts represents a probability distribution?


a)

X | P(x)

 0 | 0.1

 1 | 0.2

 2 | 0.4

 3 |  0.1


b)

 X | P(x)

 0 | 0.3

 1 | 0.2

 2 | 0.1

 3 | 0.35


c)

 X | P(x)

 0 | -0.1

 1 | 0.1

 2 | 0.4

 3 | 0.6


d)

 X | P(x)

 0 | 0

 1 | 0

 2 | 1

 3 | 0



1
Expert's answer
2022-05-03T02:57:00-0400
"P(x)\ge0, \sum _iP(x_i)=1"

a)

"0.1+0.2+0.4+0.1=0.8\not=1"

b)

"0.3+0.2+0.1+0.35=0.95\not=1"

c)

"P(0)=-0.1<0"

d)

"0+0+1+0=1"



Answer

d)

 X | P(x)

 0 | 0

 1 | 0

 2 | 1

 3 | 0